OpenStudy (anonymous):

You are performing four independent Bernoulli trials with p = 0.4 and q = 0.6. Calculate the probability of no successes.

OpenStudy (kropot72):

\[P(0\ successes)=(0.6)^{4}=you\ can\ calculate\]

OpenStudy (anonymous):

.1296? and would I only use that formula for this type of question?

OpenStudy (anonymous):

and how come the .4 isn't used?

OpenStudy (kropot72):

Your answer is correct. The probability distribution for the various possible numbers of successes, 0 to n, in n Bernoulli trials is the binomial distribution. \[\left(\begin{matrix}n \\ r\end{matrix}\right)p ^{r}q ^{n-r}\] Applying the formula to your question: \[\left(\begin{matrix}4 \\ 0\end{matrix}\right)0.4^{0}0.6^{4}=(0.6)^{4}\]

OpenStudy (anonymous):

oh okay, so that formula just deals with numbers of success?

OpenStudy (kropot72):

The formula deals with the probability of r successes and n - r failures.

OpenStudy (anonymous):

okay, thank you so much!

OpenStudy (kropot72):

You're welcome :)